# Claas Voelcker

### v1

@@ 4 Analysis of decision-aware losses in stochastic environments
While the two losses behave similarly in deterministic environments when controlling for implementation details, it is an open question if this holds in a stochastic environment. While previous work has noted that MuZero is not a good loss function in this case [6], we presented a proper theoretical treatment of the issue and an analysis of comparable algorithms. Full proofs are found in subsection A.1.
We first establish that the IterVMAL loss leads to an unbiased solution in the infinite sample limit, even when restricting the model class to deterministic functions. Note that this is a unique advantage of value-aware models, which is not true for observation-space or latent self-prediction losses such as those used in Dreamer [29] or MBPO [33].
We then show that the same is not true for MuZero, even when choosing a probabilistic model class that contains the ground-truth environment. This highlights that while the losses perform largely similar in deterministic environments, the same does not hold in the stochastic case. Finally, we verify the theoretical results empirically by presenting several stochastic extensions of environments from the DMC benchmarking suite [69].
@@ 4.1 Restricting IterVAML to deterministic models
MuZero’s formulation contains a deterministic model. This raises the question whether a deterministic model is sufficient for learning a value-equivalent model. Showing the existence of such a model relies the continuity of the transition kernel and involved functions $\phi$ and $V$. A full discussion on these conditions is presented in Appendix A, together with a proof of the existence of the deterministic model.
To reduce notational complexity, we ignore the action dependence in all following propositions, all results hold without loss of generality for the action-conditioned case as well.
@@
From this we can conclude that given sufficient samples and a sufficiently flexible function class for $\mathcal{F}$, IterVAML can recover an optimal deterministic model for value function prediction. Note that our conditions ensure the existence of a measurable model $f$, so the learning problem might still be very challenging. Nonetheless, even without the assumption that the perfect model $f$ is learnable, the IterVAML loss finds the function that is closest to it in the mean squared error sense.
4.2 Sample bias of MuZero
While MuZero is a sound loss function for deterministic models, the same is not true in stochastic environments. While alternatives to MuZero have been proposed for stochastic environments [6], to the best of our knowledge we are the first to present a formal analysis of the impact of stochasticity on the sample-based version of the MuZero loss in the population limit. While the rest of the paper focuses on continuous state-spaces, this result also holds for discrete state-spaces.
Figure 3: The source of the MuZero bias. Even though both $x_{1}$ and $\hat{x}_{1}$ are drawn from the same distribution, their value functions do not coincide.
To highlight that the problem is neither due to the $n$-step formulation nor due to suboptimal choices of model and value function class, we show that the MuZero loss does not recover the true model and value function even when they are included in the function class. Intuitively, the problem results from the fact that two sample paths drawn from the model and from the environment do not coincide, even when the true model and the learned model are equal (compare Figure 3).
@@ 4.2 Sample bias of MuZero
These results allow us to assert that the MuZero loss will not recover the correct value functions and models in environments with stochastic transitions, even when the correct model and value function are representable. On the other hand, IterVAML can recover an optimal model in stochastic environments, and will generally have an unbiased solution.
Note that the bias is dependent on the variance of the value function with regard to the transition distributions. This means that in some stochastic environments the MuZero loss might still lead perform well. But as the variance of the value function over the next state distributions increases, we expect the bias to impact the found solution.
@@ 4.3 Empirical validation of the performance in stochastic environments
Figure 4: Comparison of the impact of noise across all algorithms. While both IterVAML and MuZero are impacted by the level of noise, we observe a significantly larger drop in MuZero, which does not perform above the BYOL baseline at the highest noise level.
@@
To showcase that the bias of MuZero is not merely a mathematical curiosity, we tested the two losses in the challenging humanoid-run tasks from the DMC benchmarking suite [69] with different noise distributions applied to the action (details on the environment can be found in Appendix B). All experiments are repeated eight times and the shaded areas mark three $\sigma$ of the standard error of the mean over random seeds, which represents a $99.7\%$ certainty interval.
@@
As shown in Figure 4, we do see a clear difference between the performance of IterVAML and MuZero when increasing the noise level. At small levels of noise, all algorithms retain their performance. This is expected, since small noise on the actions can increase the robustness of a learned policy or improve exploration [32]. At large levels of noise, however, the MuZero model drops in performance to the value-unaware baseline.
It is important to note that humanoid-run is a challenging test environment and additional noise corruption does not simplify the task, which is why we do not expect IterVAML to recover its full performance in the stochastic version. Furthermore, as highlighted before, the additional stabilizing latent prediction loss is necessary for IterVAML and introduces some level of bias (for details see subsection A.3). This raises an important direction for future work, establishing a mechanism for stabilizing value-aware losses that does not introduce bias in stochastic environments.
5 Evaluating model use and environment choice
@@
Figure 5: Performance comparison overall test environments. We see that IterVAML performs slightly above MuZero in several environments, but gains from the value-awareness can only be observed in the challenging Humanoid-run environment.
@@
After investigating the impact of action noise on different loss function, we now present empirical experiments to further investigate how different implementations of $\lambda$-AC behave in empirical settings. Theoretical results [17, 16] show that value-aware losses should perform best when used in environments where learning a correct model is impossible due to access to finite samples or model capacity. Even though the expressivity of neural network models has increased in recent years, we argue that such scenarios are still highly relevant in practice. Establishing the necessary size of a model a priori is often impossible, especially since RL is deployed in scenarios where the complexity of a solution is unclear. Increasing model capacity often comes at greatly increased cost [35], which makes more efficient alternatives desirable for practitioners with limited resources. Therefore, we empirically verify under what conditions decision-aware losses show performance improvements over the simple BYOL loss.
To replace the MCTS based policy in discrete state-space MuZero, we introduced the use of model value gradients with decision-aware losses. However, all loss functions were originally designed only for improving value function estimation. We verify that they still provide benefit for policy gradient approximation by replacing the model-based policy gradient with a model-free version.
@@
For these experiments, we use three environments from the popular DMC suite [69], Walker-run, Quadruped-run, and Humanoid-run. These environments were picked from three different levels of difficulty [30] and represent a cross-section of different challenge levels in the benchmark, with Humanoid-run being a serious challenge for most established algorithms.
@@
The full algorithmic framework and all model design choices are summarized in Appendix B. The model implementation follows Hansen et al. [30]. Reported rewards are collected during training, not in an evaluation step, the same protocol used by Hansen et al. [30]. All experiments are repeated eight times and the shaded areas mark three $\sigma$ of the standard error of the mean over random seeds.
@@
The results in Figure 5 show that both MuZero and IterVAML provide a benefit over the BYOL loss in the most challenging environment, Humanoid-run. This is in line with expectations, as modelling the full state space of Humanoid-run is difficult with the model architectures we used. For smaller networks, we see a clear performance benefit in Figure 6 from the MuZero loss.
@@
Figure 7: Comparison of the impact removing model policy gradients in all environments. Overall we see decreases in performance across all environments and loss functions, with a drastic decrease in performance for IterVAML in the hardest task. Dashed lines show the mean results of the model-based gradient for comparison.
@@
As seen in Figure 7, all losses and environments benefit from better policy estimation using the model. Therefore it is advisable to use a $\lambda$-AC model both for gradient estimation and value function improvement. Compared to MuZero, IterVAML loses more performance without policy gradient computation in the hardest evaluation task. It has been noted that model-based value function estimates might be more useful for policy gradient estimation than for value function learning [5, 23]. In addition, the grounding of the MuZero loss in real rewards leads to better value function prediction in the Humanoid-run environment. Therefore a better gradient can be obtained with MuZero when the model is not used for policy gradient estimation, since a correct value function estimate becomes more important.
@@ 6 Related work
Farahmand [16] established the IterVAML formulation based on earlier work [17]. Several other extensions based on this formulation have been proposed [70, 1]. The idea of combining IterVAML with latent spaces was first developed by Abachi et al. [2], but no experimental results were provided. MuZero [61, 71] is build based on several earlier works which introduce the ideas of learning a latent model jointly with the value function [64, 53]. However, to the best of our knowledge, none of these works highlight the importance of considering the bias in stochastic environments that result from such a formulation. Antonoglou et al. [6] propose an extension to MuZero in stochastic environments, but focus on the planning procedure, not the biased loss. Hansen et al. [30] adapted the MuZero loss to continuous control environments but did not extend their formulation to stochastic variants. Grimm et al. [25] and Grimm et al. [26] consider how the set of value equivalent models evolves under choices sets of value functions. They are the first to highlight the close connection between the notions of value-awareness (or equivalence) and MuZero.
Other decision-aware algorithms
@@
Several other works propose decision-aware variants that do not minimize a value function difference. D’Oro et al. [13] weigh the samples used for model learning by their impact on the policy gradient. Nikishin et al. [52] uses implicit differentiation to obtain a loss for the model function with regard to the policy performance measure. To achieve the same goal, Eysenbach et al. [15] and Ghugare et al. [23] choose a variational formulation in the control-as-inference framework. Modhe et al. [48] proposes to compute the advantage function resulting from different models instead of using the value function. Ayoub et al. [7] presents an algorithm based on selecting models based on their ability to predict value function estimates and provide regret bounds with this algorithm.
Bisimulation and representation learning
Another line of research that focuses on similar goals as value-aware prediction is the concept of bisimulation metrics [19, 20, 72, 37, 38], behavioral similarity [4, 10] and other state-abstraction techniques [55, 56, 42, 11, 40]. While some connections between value-awareness and state abstractions and representations have been established [22, 25, 2], our work highlights that the interplay between both is vital to obtain good empirical results. Bisimulation is a stricter condition than value equivalence for a given value function, as the model needs to be equivalent with regard to the reward for all policies. Finally, we do not focus on exploration in this paper, but Guo et al. [27] show that a similar loss to ours can be exploited for targeted exploration.
Learning with suboptimal models
@@
Several works have focused on the broader goal of using models with errors without addressing the loss functions of the model. Among these, several focus on correcting models using information obtained during exploration [34, 67, 49, 58], or limiting interaction with wrong models [9, 33, 3]. Several of these techniques can be applied to improve the value function estimate of a $\lambda$-AC world model further.
@@ 7 Conclusions
In this paper, we investigated the $\lambda$-AC framework for model-based reinforcement learning with decision-aware models. The core components of all algorithms in this framework are a) a latent world model, b) a value-aware loss function and c) a model-based actor-critic approach. Empirically, we show that the design decisions established for MuZero are useful for all $\lambda$-AC algorithms. Previously established performance differences [44, 70] can be explained by this design decision. We furthermore establish a formal limitation on the performance of MuZero in stochastic environments, and verify this empirically. Finally, we conduct a series of experiments to establish which algorithmic choices lead to good performance in empirical settings.
Our results highlight the importance of decision-aware model learning in continuous control and allow us to make algorithmic recommendations. When the necessary capacity for an environment model cannot be established, using a decision-aware loss will improve the robustness of the learning algorithm with regard to the model capacity. In deterministic environments, MuZero is a good choice of loss, as the use of real rewards provides a better learning signal for the value function. In stochastic environments, an IterVAML-based design is more effective, as the IterVAML loss is unbiased.
Overall, we find that $\lambda$-AC is an important addition to the RL toolbox in complex environments where other modelling approaches fail. In future work, evaluating other design decisions, such as probabilistic models [23] and alternative exploration strategies [30, 27] can provide important insights
Acknowledgments and Disclosure of Funding
We would like to thank the members of the Adaptive Agents Lab, the Toronto Intelligent Systems Lab, David Glukhov, and Animesh Garg, who provided feedback on the ideas, the draft of this paper, and provided additional ressources for running all experiments. AMF acknowledges the funding from the Canada CIFAR AI Chairs program, as well as the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. CV and RA thank Brandon Amos for providing helpful feedback at an early stage of the project.
Appendix A Proofs

### v3

@@ 4 Analysis of decision-aware losses in stochastic environments
After accounting for model implementation, both MuZero and IterVAML behave similarly in experiments. However, it is an open question if this holds in general. Almost all commonly used benchmark environments in RL are deterministic, yet stochastic environments are an important class of problems both for theoretical research and practical applications. Therefore, we first present a theoretical investigation into the behavior of both the MuZero and IterVAML losses in stochastic environments, and then evaluate our findings empirically. Proofs are found in Subsection A.1.
In Proposition 1 we establish that $\lambda$-IterVAML leads to an unbiased solution in the infinite sample limit, even when restricting the model class to deterministic functions under measure-theoretic conditions on the function class. This is a unique advantage of value-aware models. Algorithms such as Dreamer (Hafner et al., 2020) or MBPO (Janner et al., 2019) require stochastic models, because they model the complete next-state distribution, instead of the expectation over the next state.
We then show in Proposition 2 that MuZero’s joint model- and value function learning algorithm leads to a biased solution, even when choosing a probabilistic model class that contains the ground-truth environment. This highlights that while the losses are similar in deterministic environments, the same does not hold in the stochastic case. Finally, we verify the theoretical results empirically by presenting stochastic extensions of environments from the DMC suite (Tunyasuvunakool et al., 2020).
@@ 4.1 Restricting IterVAML to deterministic models
In most cases, deterministic function approximations cannot capture the transition distribution on stochastic environments. However, it is an open question whether a deterministic model is sufficient for learning a value-equivalent model, as conjectured by (Oh et al., 2017). We answer this now in the affirmative. Showing the existence of such a model relies on the continuity of the transition kernel and involved functions $\phi$ and $V$ [^2] To reduce notational complexity, we ignore the action dependence in all following propositions; all results hold without loss of generality for the action-conditioned case as well.
@@
Furthermore, the same $f^{*}$ is a minimizer of the expected IterVAML loss:
@@
We can conclude that given a sufficiently flexible function class $\mathcal{F}$, IterVAML can recover an optimal deterministic model for value function prediction. Note that our conditions solely ensure the existence of a measurable function; the learning problem might still be very challenging. Nonetheless, even without the assumption that the perfect model $f$ is learnable, the IterVAML loss finds the function that is closest to it in the mean squared error sense, as shown by Farahmand (2018).
4.2 Sample bias of MuZero’ value function learning algorithm
MuZero is a sound algorithm for deterministic environments, however, the same is not true for stochastic ones. While changed model architecture for MuZero have been proposed for stochastic cases (Antonoglou et al., 2022), the value function learning component contains its own bias. To highlight that the problem is neither due to the $n$-step formulation nor due to suboptimal architecture or value function class, we show that the value function estimate does not converge to the model Bellman target using the MuZero loss. Intuitively, the problem results from the fact that two samples drawn from the model and from the environment do not coincide, even when the true model and the learned model are equal (see Figure 3). This bias is similar to the double sampling issue in Bellman residual minimization, but is distinct as the introduction of the stop-gradient in MuZero’s loss function does not address the problem.
Figure 3: The source of the MuZero bias. Though $x_{1}$ and $\hat{x}_{1}$ are drawn from the same distribution, their values do not coincide.
@@ 4.2 Sample bias of MuZero’ value function learning algorithm
Assume a non-deterministic MDP with a fixed, but arbitrary policy $\pi$, and let $p$ be the transition kernel. Let $\mathcal{V}$ be an open set of functions, and assume that it is Bellman complete: $\forall V\in\mathcal{V}:\mathcal{T}V\in\mathcal{V}$.
@@
The bias indicates that MuZero will not recover the correct value function in environments with stochastic transitions, even when the correct model is used and the function class is Bellman complete. On the other hand, model-based MVE such as used in $\lambda$-IterVAML can recover the model’s value function in stochastic environments.
The bias is dependent on the variance of the value function with regard to the transition distributions. This means that in some stochastic environments the MuZero loss might still perform well. But as the variance of the value function increases, the bias to impact the solution.
If the MuZero loss is solely used for model learning and the value function is learned fully model-free or model-based, the IterVAML and MuZero algorithms show strong similarities (compare Subsection A.2). The main difference is that MuZero uses a bootstrap estimate for the value function, while IterVAML uses the value function estimate directly. However, when jointly training value function and model in a stochastic environment, neither the model nor the value function converge to the correct solution, due to the tied updates.
@@ 4.3 Empirical validation of the performance in stochastic environments
Figure 4: Comparison of the impact of noise across all algorithms. While both IterVAML and MuZero are impacted by the level of noise, we observe a larger drop in MuZero, which does not perform above the BYOL baseline at the highest noise level.
@@
To showcase that the bias of MuZero’s value learning strategy is not merely a mathematical curiosity, we tested the two losses in the challenging humanoid-run tasks from the DMC benchmarking suite (Tunyasuvunakool et al., 2020) with different noise distributions applied to the action (details on the environment can be found in Appendix C).
@@
As shown in Figure 4, we do see a clear difference between the performance of $\lambda$-IterVAML and $\lambda$-MuZero when increasing the noise level. At small levels of noise, all algorithms retain their performance. This is expected, since small noise on the actions can increase the robustness of a learned policy or improve exploration (Hollenstein et al., 2022). At large levels of noise, however, the $\lambda$-MuZero algorithm drops in performance to the value-agnostic baseline. We provide additional experiments with more stochastic environments in Appendix D which further substantiate our claims.
@@
Figure 5: Performance comparison overall test environments. We see that IterVAML performs slightly above MuZero in several environments, but decisive gains from the value-aware losses ($\lambda$-IterVAML, $\lambda$-MuZero) over BYOL can only be observed in the challenging Humanoid-run environment.
It is important to note that humanoid-run is a challenging environment and strong noise corruption increases difficulty, therefore $\lambda$-IterVAML will not retain the full performance in the stochastic version. Furthermore, as highlighted before, the stabilizing latent prediction loss is necessary for $\lambda$-IterVAML and introduces some level of bias (for details see Subsection A.4). However, as all algorithms are impacted by the stabilization term, it is still noticeable that $\lambda$-MuZero’s performance drops more sharply. This raises an important direction for future work, establishing a mechanism for stabilizing value-aware losses that does not introduce bias in stochastic environments.
5 Evaluating model capacity and environment choice
@@
After investigating the impact of action noise on different loss function, we now present empirical experiments to further investigate how different implementations of $\lambda$ algorithms behave in empirical settings. Theoretical results (Farahmand et al., 2017; Farahmand, 2018) show that value-aware losses perform best when learning a correct model is impossible due to access to finite samples or model capacity. Even though the expressivity of neural networks has increased in recent years, we argue that such scenarios are still highly relevant in practice. Establishing the necessary size of a model a priori is often impossible, since RL is deployed in scenarios where the complexity of a solution is unclear. Increasing model capacity often comes at greatly increased cost (Kaplan et al., 2020), which makes more efficient alternatives desirable. Therefore, we empirically verify under what conditions decision-aware losses show performance improvements over the simple BYOL loss.
@@
For these experiments, we use environments from the popular DMC suite (Tunyasuvunakool et al., 2020), Walker-run, Quadruped-run, and Humanoid-run. These environments were picked from three different levels of difficulty (Hansen et al., 2022) and represent different challenge levels in the benchmark, with Humanoid-run being a serious challenge for most established algorithms.
@@
The algorithmic framework and all model design choices are summarized in Appendix C. The model implementation follows Hansen et al. (2022). Reported rewards are collected during training, not in an evaluation step, the same protocol used by Hansen et al. (2022).
@@
The results in Figure 5 show that both MuZero and IterVAML provide a benefit over the BYOL loss in the most challenging environment, Humanoid-run. This is expected, as modelling the full state space of Humanoid-run is difficult with the model architectures we used. For smaller networks, we see a stable performance benefit in Figure 6 from the MuZero loss. $\lambda$-IterVAML also outperforms the BYOL baseline, but fails to achieve stable returns.
@@
Figure 7: Comparison of the impact removing model policy gradients in all environments. We see decreases in performance across all environments and loss functions, with a drastic decrease in performance for IterVAML in humanoid. Dashed lines show the mean results of the model-based gradient for comparison.
@@
As seen in Figure 7, all losses and environments benefit from better policy estimation using the model. Therefore it is advisable to use a $\lambda$ model both for gradient estimation and value function improvement. Compared to MuZero, IterVAML loses more performance without policy gradient computation in the hardest evaluation task. It has been noted that model-based value function estimates might be more useful for policy gradient estimation than for value function learning (Amos et al., 2021; Ghugare et al., 2023). In addition, the grounding of the MuZero loss in real rewards likely leads to better value function prediction in the Humanoid-run environment. Therefore a better update can be obtained with MuZero when the model is not used for policy gradient estimation, since learning is driven by the value function update.
@@ 6 Related work
VAML and MuZero:  Farahmand (2018) established IterVAML based on earlier work (Farahmand et al., 2017). Several extensions to this formulation have been proposed, such as a VAML-regularized MSE loss (Voelcker et al., 2022) and a policy-gradient aware loss (Abachi et al., 2020). Combining IterVAML with latent spaces was first developed by Abachi et al. (2022), but no experimental results were provided. MuZero (Schrittwieser et al., 2020; Ye et al., 2021) is build based on earlier works which introduce the ideas of learning a latent model jointly with the value function (Silver et al., 2017; Oh et al., 2017). However, none of these works highlight the importance of considering the bias in stochastic environments that result from such a formulation. Antonoglou et al. (2022) propose an extension to MuZero in stochastic environments, but focus on the planning procedure, not the biased value function loss. Hansen et al. (2022) adapted the MuZero loss to continuous control environments but decisiondid not extend their formulation to stochastic variants. Grimm et al. (2020) and Grimm et al. (2021) consider how the set of value equivalent models relates to value functions. They are the first to highlight the close connection between the notions of value-awareness and MuZero.
@@
Other decision-aware algorithms:  Several other works propose decision-aware variants that do not minimize a value function difference. D’Oro et al. (2020) weigh the samples used for model learning by their impact on the policy gradient. Nikishin et al. (2021) uses implicit differentiation to obtain a loss for the model function with regard to the policy performance measure. To achieve the same goal, Eysenbach et al. (2022) and Ghugare et al. (2023) choose a variational formulation. Modhe et al. (2021) proposes to compute the advantage function resulting from different models instead of using the value function. Ayoub et al. (2020) presents an algorithm based on selecting models based on their ability to predict value function estimates and provide regret bounds with this algorithm.
@@
Learning with suboptimal models:  Several works have focused on the broader goal of using models with errors without addressing the loss functions of the model. Among these, several focus on correcting models using information obtained during exploration (Joseph et al., 2013; Talvitie, 2017; Modi et al., 2020; Rakhsha et al., 2022), or limiting interaction with wrong models (Buckman et al., 2018; Janner et al., 2019; Abbas et al., 2020). Several of these techniques can be applied to improve the value function of a $\lambda$ world model further. Finally, we do not focus on exploration, but Guo et al. (2022) show that a similar loss to ours can be exploited for targeted exploration.
@@ 7 Conclusions
In this paper, we investigated model-based reinforcement learning with decision-aware models with three main question focused on (a) implementation of the model, (b) theoretical and practical differences between major approaches in the field, and (c) scenarios in which decision-aware losses help in model-based reinforcement learning.
Empirically, we show that the design decisions established for MuZero are a strong foundation for decision-aware losses. Previous performance differences (Lovatto et al., 2020; Voelcker et al., 2022) can be overcome with latent model architectures. We furthermore establish a formal limitation on the performance of MuZero in stochastic environments, and verify this empirically. Finally, we conduct a series of experiments to establish which algorithmic choices lead to good performance empirically.
Our results highlight the importance of decision-aware model learning in continuous control and allow us to make algorithmic recommendations. When the necessary capacity for an environment model cannot be established, using a decision-aware loss will improve the robustness of the learning algorithm with regard to the model capacity. In deterministic environments with deterministic models, MuZero’s value learning approach can be a good choice, as the use of real rewards seemingly provides a grounded learning signal for the value function. In stochastic environments, a model-based bootstrap is more effective, as the model-based loss does not suffer from MuZero’s bias.
Overall, we find that decision aware learning is an important addition to the RL toolbox in complex environments where other modelling approaches fail. However, previously established algorithms contain previously unknown flaws or improper design decisions that have made their adoption difficult, which we overcome. In future work, evaluating other design decisions, such as probabilistic models (Ghugare et al., 2023) and alternative exploration strategies (Hansen et al., 2022; Guo et al., 2022) can provide important insights.
Appendix A Proofs and mathematical clarifications